three candidates are contesting in an election. Candidates a and b have same chance of wining but candidates c has twice the chance of b and a find
1:probability that a wins
2:probability that c does not win
1
Expert's answer
2012-01-26T12:12:09-0500
Let p be the probability of winning for a.
By assumption it is also equal to
the probability of winning of b.
Moreover, the probability of winning for c
is twice the chance of b and a, and so it is equal to 2p.
Since only one
candidate can win, we have
p + p + 2p = 1
4p=1

p=1/4=0.25

Hence
1) the probability that a win is 0.25
2) the
probability that c does not win is 1-2p=0.5